### v1

# Extraction of classical ergotropy

Abstract
Finding the time dependent perturbation that extracts the maximal amount of energy (a.k.a. ergotropy) from a thermally isolated quantum system is a central, solved, problem in quantum thermodynamics. Notably, the same problem has been long studied for classical systems as well, e.g., in the field of plasma physics, but a general solution is still missing there. By building on the analogy with the quantum solution, we provide the classical ergotropy extraction driving: it consists of an instantaneous quench followed by an adiabatic return. We illustrate how the solution, which is valid under an ergodic assumption, is instrumental to finding the ergotropy extracting driving in more general cases. We also show that, just like in the quantum case, the classical ergotropy splits into a coherent and an incoherent part. The presented results open new ways for practical energy recovery in the classical regime while suggesting that there is nothing genuinely quantum in the quantum ergotropy problem.
What is the maximal amount of energy that can be extracted from a thermally isolated system by means of a cyclical external driving? And what is the driving that achieves such maximal energy extraction?
These questions are in the limelight of current investigation in the field of quantum thermodynamics [1, 2, 3, 4] where they naturally emerge in the study of quantum batteries [5, 6, 7, 8], quantum heat engines [9, 10, 11, 12, 13] and dynamic cooling [14, 15, 16, 17]. While addressing these questions for quantum systems has a long history [18] they took momentum only with the work of Allahverdyan et al. [19], who coined the expression “ergotropy” for the maximal energy extractable from a quantum thermally isolated system, and provided its analytical expression along with the expression of the according "ergotropy extracting" unitary operator, see Eqs. (9,13) below.
Most notably the ergotropy extraction problem has been long studied for classical systems as well. For example, in the field of plasma physics, where, e.g., the recovery of the maximal amount of energy from a plasma of fusion products constitutes a pressing practical challenge [20, 21, 22, 23, 24].
In the pioneering and elegantly succinct work of Ref. [25] Gardner pinpointed the fundamental principles that guide the calculation of the maximal energy extraction from a classical system by means of phase space volume preserving maps, a.k.a. "Gardner free energy" [20, 21]. As pointed out recently [23] any volume preserving map can be arbitrarily well approximated by a symplectic map, therefore, formally, the Gardner free energy is indeed the maximal energy that can be extracted by a Hamiltonian flow, i.e., the classical ergotropy. Therefore, as of today, determining the classical ergotropy is a problem with a known solution.
However, the pressing practical question of what is the classical ergotropy extraction driving remains an open question. Here we first re-derive the classical ergotropy and express it in a new way, Eq. (7), and then demonstrate a general driving protocol that achieves it. The protocol is intuituve, physically motivated and is inspired by the according solution of the quantum problem whose close analogy with the classical problem is demonstrated. The presented protocol works under the provision of the ergodic hypothesis, however we shall discuss how it in fact can be employed as the basis for designing the ergotropy extraction drivings in more general cases as well. This opens new ways for the practical implementation of energy extraction protocols in the classical regime. The presented theory also sheds further light on the quantum ergotropy, showing that there is very little (if not nothing at all) genuinely quantum in it, besides the discreteness of variables.
Gardner free energy (a.k.a. classical ergotropy).– Given a classical system with unperturbed Hamiltonian $H_{0}(\mathbf{z})$ being initially in a statistical state described by the phase space distribution $\rho_{0}(\mathbf{z})$, Gardner asked what is the distribution $\rho_{1}(\mathbf{z})$ featuring the smallest energy expectation among all those that can be connected to $\rho_{0}$ by a phase-volume preserving map [25]. Gardner concluded that (i) $\rho_{1}(\mathbf{z})$ must be a decreasing function $g$ of the system unperturbed Hamiltonian $H_{0}(\mathbf{z})$, (ii) for any positive real number $\sigma$, the volume of phase space where $\rho_{1}>\sigma$ must be equal to that where $\rho_{0}>\sigma$  [^1] A simple aftermath of Gardner’s prescription is that one cannot extract energy from states of the form $\rho_{0}=g(H_{0})$ with $g$ a decreasing fuction, a fact that has been formally proved also in Ref. [27], see also [28] for a more accessible proof.. These prescriptions form the basis of the so-called "Gardner restacking algorithm" and allow to calculate the "Gardner free energy" and the function $g$ that determines the “Gardner ground state” $\rho_{1}=g(H_{0})$ [20, 21], a.k.a. the “passive state” relative to $H_{0}$ [27, 28]. Let
$R(\sigma)=\int d\mathbf{z}\theta[\rho_{0}(\mathbf{z})-\sigma]$ (1)
denote the measure of phase space space where $\rho_{0}>\sigma$ (the symbol $\theta[\cdot]$ denotes Heaviside step function). $R$ is a decreasing, hence invertible, function. Let
$\Omega_{0}(E)=\int d\mathbf{z}\theta[E-H_{0}(\mathbf{z})]$ (2)
[…]
$R(\sigma)=\int d\mathbf{z}\theta[g(H_{0}(\mathbf{z}))-\sigma]\,$, (3)
or, equivalently,
$R(\sigma)$ $=\int d\mathbf{z}\theta[g^{-1}(\sigma)-H_{0}(\mathbf{z})]=\Omega_{0}(g^{-1}(\sigma))\,.$ (4)
Applying $E_{0}$ to both sides of the above equation, we get
$E_{0}(R(\sigma))=g^{-1}(\sigma)\,.$ (5)
The Gardner ground state is therefore $\rho_{1}(\mathbf{z})=g(H_{0}(\mathbf{z}))$ with $g$ obtained by inverting Eq. (5)  [^2] Note that since $E_{0}$ is increasing, and $R$ is decreasing, $g^{-1}$ is decreasing, therefore $g$ is decreasing as expected.. The minimal energy reachable by volume preserving maps reads, accordingly:
$\breve{E}_{c}$ $=\int d\mathbf{z}H_{0}(\mathbf{z})g(H_{0}(\mathbf{z}))=\int_{0}^{\infty}de\omega_{0}(e)eg(e)$
$=$ $\int_{0}^{\infty}d\Phi E_{0}(\Phi)g(E_{0}(\Phi))=\int_{0}^{\infty}d\Phi E_{0}(\Phi)R^{-1}(\Phi)\,$, (6)
where in the first line we changed the integration variable from phase-space $\mathbf{z}$ to energy $e$, using the density of states $\omega_{0}(e)=\Omega_{0}^{\prime}(e)$, in the second line we first made the change of variable $e\rightarrow\Phi=\Omega_{0}(e)$, and then we used Eq. (5). Remarkably, by inverting the functions $R,\Omega_{0}$ in Eqs. (1,2) one swiftly gets $\breve{E}_{c}$, hence the Gardner free energy, that is, the classical ergotropy,
$\mathcal{E}_{c}=\int d\mathbf{z}H_{0}(\mathbf{z})\rho_{0}(\mathbf{z})-\int_{0}^{\infty}d\Phi E_{0}(\Phi)R^{-1}(\Phi)\,.$ (7)
Despite the interest of this and similar equivalent expressions [20, 21], they tell nothing about the pressing practical question of how to engineer a time dependent driving $V(\mathbf{z},t)$ such that the system state $\rho_{0}$ will indeed evolve onto the “Gardner ground state” $\rho_{1}$ (a.k.a the passive state of $\rho_{0}$) under the flow generated by the total Hamiltonian $H(\mathbf{z},t)=H_{0}(\mathbf{z})+V(\mathbf{z},t)$ in some time span $0,\tau]$.
The Quench-Adiabat protocol.– To get a handle on the above issue let’s look at the analogous quantum problem. Given a quantum system with Hamiltonian $\hat{H}_{0}$ being in a state described by the density operator $\hat{\rho}_{0}$, we ask what is the unitary evolution operator $\hat{U}$ that leads to the state $\hat{\rho}_{1}=\hat{U}\hat{\rho}_{0}\hat{U}^{\dagger}$ of lowest energy expectation
$\breve{E}_{q}=\min_{\hat{U}}\text{Tr}\,\hat{H}_{0}\hat{U}\hat{\rho}_{0}\hat{U}^{\dagger}\,.$ (8)
For Hamiltonians with discrete spectrum acting on a Hilbert space of finite dimension $d$ the above minimum is reached by the unitary operator [19]
$\hat{U}=\sum_{k=1}^{d}|e_{k}\rangle\langle r_{k}|\,$, (9)
[…]
Under the evolution $\hat{U}$ the system reaches the minimal energy state
$|\displaystyle\hat{\rho}_{1}=\hat{U}\hat{\rho}_{0}\hat{U}^{\dagger}=\sum_{k}r_{k}\ket{e_{k}}\bra{e_{k}}$ (12)
and the quantum ergotropy, i.e., the difference between the initial energy expectation $E_{0}=\text{Tr}\,\hat{H}_{0}\hat{\rho}_{0}$ and $\breve{E}_{q}$ amounts to
$\mathcal{E}_{q}=\sum_{j,k}r_{j}(|\langle r_{j}|e_{k}\rangle|^{2}-\delta_{jk})e_{k}\,$, (13)
where $\delta_{jk}$ denotes the Kronecker delta.
Note that the optimal unitary $\hat{U}$ is exactly realising the quantum analogue of Gardner prescriptions [25]: the “ground state” $\hat{\rho}_{1}$, Eq. (12) is a decreasing function of $\hat{H}_{0}$, and the number of eigenstates of initial and final density operators ($\hat{\rho}_{0}$ and $\hat{\rho}_{1}$), whose eigenvalue is less than some value $\sigma$ are the same, for any $\sigma>0$. This is so because the density operator eigenvalues are invariant under unitary evolution. In classical mechanics, the analogous condition is dictated by the fact that the phase space density is invariant under Liouville evolution.
What time-dependent perturbation $\hat{V}(t)$ should one enact so that the according Hamiltonian
$\hat{H}(t)=\hat{H}_{0}+\hat{V}(t)$ (14)
would induce the dynamical evolution $\hat{U}$? Notably the answer is not unique. From the mathematical point of view, perhaps the simplest solution is to chose $\hat{H}(t)=\hat{K}\doteq\hbar\frac{i}{\tau}\log U$, for $t\in[0,\tau]$, so that trivially one has $e^{-i\hat{K}\tau/\hbar}=\hat{U}$ in the time span $0,\tau]$. From the physical viewpoint, we note that $\hat{U}$ can be implemented by the following quench-adiabat (QA) protocol: (i) instantaneously quench at $t=0$ to
$\hat{H}_{1}=f(\hat{\rho}_{0})$ (15)
[…]
In classical mechanics one can do exactly the same QA protocol: (i) instantaneously quench to
$H_{1}(\mathbf{z})=f(\rho_{0}(\mathbf{z}))$ (16)
[…]
To prove that the QA protocol extracts the calssical ergotropy let
$P_{1}(\Omega)$ $=\int d\mathbf{z}\rho_{0}(\mathbf{z})\delta[\Omega-\Omega_{1}(H_{1}(\mathbf{z}))]$ (17)
denote the probability density of finding, at $t=0$, the system on the iso-$H_{1}$ hyper-surface that encloses the volume of phase space $\Omega$. The symbol $\delta$ denotes Dirac’s delta and
$\Omega_{1}(E)=\int d\mathbf{z}\theta(E-H_{1}(\mathbf{z}))$ (18)
denotes the volume of phase space enclosed by the hypersurface $H_{1}=E$. We have
$P_{1}(\Omega)$ $=\int d\mathbf{z}f^{-1}(H_{1}(\mathbf{z}))\delta[\Omega-\Omega_{1}(H_{1}(\mathbf{z}))]$
$=\int de\omega_{1}(e)f^{-1}(e)\delta[\Omega-\Omega_{1}(e)]$
$=\int d\Phi f^{-1}(E_{1}(\Phi))\delta[\Omega-\Phi]=f^{-1}(E_{1}(\Omega)),$ (19)
where in the first line we used $\rho_{0}=f^{-1}(H_{1})$, in the second line we changed the integration variable from phase-space $\mathbf{z}$ to energy $e$, using the density of states $\omega_{1}(e)=\Omega_{1}^{\prime}(e)$, and in the third line we made the change of variable $e\rightarrow\Phi=\Omega_{1}(e)$, with $E_{1}=\Omega_{1}^{-1}$ denoting its inverse.
Under the ergodic hypothesis, because of adiabatic invariance of the phase volume, the quantity $P_{1}(\Omega)$ is in fact equal to the probability density of finding the system on a hypersurface of constant $H_{0}$, enclosing a volume $\Omega$, at the end of the adiabatic driving [42, 28]. Therefore the final mean energy reads:
$\breve{E}_{c}=\int d\Omega P_{1}(\Omega)E_{0}(\Omega)\,.$ (20)
As shown in the appendix, the above expression is in fact identical to Gardner ground state energy Eq. (7), thus concluding the proof.
Quantum-classical analogy.– The classical ergotropy amounts to $\mathcal{E}_{c}=\int d\mathbf{z}\rho(\mathbf{z})H_{0}(\mathbf{z})-\int d\Omega P_{1}(\Omega)E_{0}(\Omega)$, or, in more symmetric form:
$\mathcal{E}_{c}=\int d\Omega[P_{0}(\Omega)-P_{1}(\Omega)]E_{0}(\Omega)\,$, (21)
where
$P_{0}(\Omega)=\int d\mathbf{z}\delta[\Omega-\Omega_{0}(H_{0}(\mathbf{z}))]\rho_{0}(\mathbf{z})$ (22)
[…]
$P_{0}(\Omega)$
$=\int d\Theta\int d\mathbf{z}\delta[\Omega-\Omega_{0}(H_{0}(\mathbf{z}))]\delta[\Theta-\Omega_{1}(H_{1}(\mathbf{z}))]f^{-1}(H_{1}(\mathbf{z}))$
$=\int d\Theta\int d\mathbf{z}\delta[\Omega-\Omega_{0}(H_{0}(\mathbf{z}))]\delta[\Theta-\Omega_{1}(H_{1}(\mathbf{z}))]f^{-1}(E_{1}(\Theta))$
$=\int d\Theta G[\Theta|\Omega]P_{1}(\Theta)\,$, (23)
where we first inserted the unit resolution, $1=\int d\Theta\delta(\Theta-y)$, and then we introduced the “overlap” between the microcanonical states of fixed $H_{0}(\mathbf{z})$ and $H_{1}(\mathbf{z})$, whose support respectively enclose the volumes $\Theta$ and $\Omega$  [^3] When parametrised via the enclosed volume variable $\Phi$, instead of the usual energy $E$, the microcanonical ensemble is naturally normalised and simply reads: $\rho(\mathbf{z};\Phi)=\delta(\Phi-\Omega(H(\mathbf{z})))$, where $\Omega(x)$ is the phase space volume of the region $H\leq x$ [28]. $G[\Theta|\Phi]$ can be interpreted as the conditional probability that the system is found on the hypersurface of constant $H_{0}$ that encloses the volume $\Omega$ given that it is homogeneously (i.e., microcanonically) distributed on the hypersurface of constant $H_{1}$ that encloses the volume $\Theta$, or viceversa. It is a special instance of the propagator $G_{0,\tau}[\Theta|\Omega]=\int d\mathbf{z}\delta[\Omega-\Omega_{0}(H_{0}(\mathbf{z}))]\delta[\Theta-\Omega_{1}(H_{1}(\mathbf{z}_{\tau}))]$ valid for a generic time dependent protocol taking $H_{0}$ to $H_{1}$ in a finite time $\tau$, where $\mathbf{z}_{\tau}$ denotes the evolved of initial condition $\mathbf{z}$ according to the protocol [41, 42, 28]. For a sudden quench it is $\tau=0$, so that $\mathbf{z}_{\tau}=\mathbf{z}$ and the expression boils down to Eq. (24). Furthermore just like the quantum transition probabilities $p_{ij}$ form a doubly stochastic matrix, similarly the propagator $G$ is doubly stochastic in the sense that $\int d\Theta G[\Theta|\Omega]=\int d\Omega G[\Theta|\Omega]=1$ [42, 28]
$G[\Theta|\Omega]=\int d\mathbf{z}\delta[\Omega-\Omega_{0}(H_{0}(\mathbf{z}))]\delta[\Theta-\Omega_{1}(H_{1}(\mathbf{z}))]$ (24)
Summing up, the classical ergotropy reads:
$\mathcal{E}=\int d\Omega d\Phi P_{1}(\Theta)(G[\Theta|\Omega]-\delta[\Omega-\Theta])E_{0}(\Omega)\,.$ (25)
Note the formal analogy with the quantum formula (13) whereby the discrete (adiabatically invariant) principal quantum numbers $k,j$ are replaced by the continuous (adiabatically invariant) "enclosed volumes" variables $\Theta,\Omega$, and the quantum overlap $p_{kj}=|\langle r_{j}|e_{k}\rangle|^{2}$ is replaced by the classical overlap $G[\Theta|\Omega]$  [^4] The analogy becomes even more evident if one notes that, for systems admitting a phase space description, the quantum transitions $p_{kj}$ can be written as the phase space overlaps between the Wigner functions of the eigenstates of $\hat{H}_{0}$ and $\hat{H}_{1}$ [49] and that (for ergodic systems), such Wigner functions approach the classical microcanonical distributions in the classical limit..
Quite remarkably, just like the quantum ergotropy, the classical ergotropy splits into a coherent and an incoherent part [45]. Let
$\mathcal{D}[\rho](\mathbf{z})=\int d\mathbf{z}^{\prime}\,\rho(\mathbf{z}^{\prime})\delta[\Omega_{0}(H_{0}(\mathbf{z}^{\prime}))-\Omega_{0}(H_{0}(\mathbf{z}))]$ (26)
[…]
$\beta\mathcal{E}_{c}^{c}[\rho]$ $=C[\rho]+D[\mathcal{P}[\mathcal{D}[\rho]||\rho_{\beta}]-D[\mathcal{P}[\rho]||\rho_{\beta}]\,$, (27)
where
$\mathcal{E}_{c}^{c}[\rho]\doteq\mathcal{E}_{c}[\rho]-\mathcal{E}_{c}^{i}[\rho]\quad\mathcal{E}_{c}^{i}[\rho]\doteq\mathcal{E}_{c}[\mathcal{D}[\rho]]$ (28)
define the coherent and incoherent parts of $\mathcal{E}_{c}$, $\mathcal{P}[\rho]$ denotes the passive state (i.e., the Gardner ground state) associated to $\rho$, $\rho_{\beta}=e^{-\beta H_{0}}/Z$ is a thermal state, $D[\cdot||\cdot]$ denotes the Kullback-Leibler divergence, and
$C[\rho]=D[\rho||\mathcal{D}[\rho]]$ (29)
quantifies the “coherence” of $\rho$, that is how much it is distinguisgable from its dephased companion $\mathcal{D}[\rho]$. Equations (27-29) read exactly like their quantum counterparts where $D$ denotes the quantum Kullback-Leibler divergence, and $\mathcal{D}$ is the quantum dephasing operator relative to $\hat{H}_{0}$ [45].
Example.–
Figure 1: Panels a-d): Phase space sketch of extraction of ergotropy from a non-stationary Gaussian state, Eq. (31) of a harmonic oscillator, Eq. (30). a) The phase density rotates under the unperturbed dynamics. b) At $t=0$ the harmonic potential gets instantaneously displaced so as to lock the state. c,d) In the time span $0,\tau\gg 1]$ the potential is slowly returned to its initial position. Panels e-f): Sketches of the states $\rho_{0}$, $\rho^{\prime}$, Eqs. (34,35). In all panels darker blue denotes higher density. The solid lines denote iso-$H_{0}$ curves in phase space
Consider a 1D Harmonic oscillator
$H_{0}(q,p)=\frac{p^{2}}{2}+\frac{q^{2}}{2}$ (30)
being at $t=0$ in the Gaussian state
$\rho_{0}(q,p)=\frac{1}{2\pi\sigma}\exp\left[-\frac{(q-q_{0})^{2}+p^{2}}{2\sigma}\right]\,.$ (31)
See the sketch in Fig. 1a). The state is clearly non-stationary (because $\{H_{0},\rho_{0}\}\neq 0$) and would evolve by rotating around the phase space origin at angular frequency $\omega=1$, without changing shape, if unperturbed. By suddenly displacing the harmonic potential minimum at $q=q_{0}$ blocks the state at the bottom of the displaced harmonic potential: the state is now stationary (in fact passive) relative to the new Hamiltonian
$H_{1}(q,p)=\frac{p^{2}}{2}+\frac{(q-q_{0})^{2}}{2}=f(\rho_{0}(\mathbf{z}))\,$, (32)
where $f(x)=-\sigma\ln(2\pi\sigma x)$ is a decreasing function, see Fig. 1b). By adiabatically moving the harmonic potential minimum back to the origin of the $q$-axis, the state evolves onto
${\rho_{1}}(q,p)=\frac{1}{2\pi\sigma}\exp\left[-\frac{(q^{2}+p^{2})}{2\sigma}\right]=\frac{e^{-H_{0}(q,p)/\sigma}}{2\pi\sigma}\,$, (33)
because each instantaneous Hamiltonian visited by the protocol is ergodic, see Fig. 1c,d). Since the state is passive relative to $H_{0}$ no more energy can be extracted therefrom.
Discussion.– There are a number of issues that limit the applicability of the QA protocol for ergotropy extraction. Let’s consider, for example, an initial state of the harmonic oscillator featuring a central valley, e.g.:
$\rho_{0}(q,p)$ $=\exp{\left[-\frac{(p^{2}+(q-q_{0})^{2})^{2}}{4}+\frac{p^{2}+(q-q_{0})^{2}}{2}\right]}/N\,$, (34)
where $N$ is the normalisation factor, see the sketch in Fig. 1e). Any choice of decrasing function $f$ would result in a Mexican hat shaped Hamiltonian $H_{1}$ which could be very hard to be implemented. Furthermore since any such $H_{1}$ would feature a central peak, not all regions $H_{1}<E$ would be simply connected, unlike those for which $H_{0}<E$, therefore they cannot be smoothly and adiabatically linked to each other. How to get around such problems? In this case one could still apply the QA protocol, but with the physical $H_{1}$ in Eq. (32), which has the same level hypersurfaces as $\rho_{0}$. Having $H_{1}$ the same topology of phase volumes as $H_{0}$, by adiabatically return to $H_{0}$, the state would be mapped onto
$\rho^{\prime}(q,p)$ $=\exp{\left[-\frac{(p^{2}+q^{2})^{2}}{4}+\frac{p^{2}+q^{2}}{2}\right]}/N=\frac{e^{-H^{2}_{0}+H_{0}}}{N}\,$, (35)
which is less energetic than $\rho_{0}$, but not yet passive although it is stationary (because it is a function of $H_{0}$, but not a decreasing one). In fact $\rho^{\prime}$ has a central valley, see the sketch in Fig. 1f). Using the quantum jargoon, one would say that the state $\rho^{\prime}$ features population inversions. In order to decrease the energy further one should permute rings of higher populations with rings of same volume with lower population and higher (average) energy. The smaller the rings thickness, the closer one can get to the ergotropy extraction, by increasing the number of such permutations. Interestingly, physical drivings that indeed achieve such permutations have been described in Refs. [47]. Notably, such permutation protocols defy the conditions of validity of the adiabatic theorem (because they aim at adiabatically connecting topologically inequivalent regions), but, as numerically demonstrated in Ref. [48], they can nonetheless approach the wanted permutation with very high accuracy. Thus the breakdown of the ergodic hypothesis does not seem to be such a dramatic problem for ergotropy extraction (as much as it is not dramatic for the mechanical foundations of thermodynamics [35, 28]).
In general it is expected that the QA protocol works well whenever $H_{0}$ and $\rho_{0}$ have topologically equivalent phase space structure. In particular no problems occur when the regions $H_{0}\leq E$ and $\rho_{0}>\sigma$ are all simply connected as is the case in the first example above.
Conclusions.– By bridging between quantum thermodynamics and classical plasma physics, the present work offers new viewpoints to both fields. On one hand it provides a guiding principle for energy extraction from classical systems based on the QA protocol (or variants thereof, e.g., as discussed above); on the other it suggests that the quantum theory of ergotropy has no genuinely quantum feature. The two pictures, classical and quantum, are fully one the close analogue of the other. Specifically a decomposition into coherent and incoherent part exists just like in the quantum case, corroborating the idea that energy coherences do not constitute a genuine quantum thermodynamic advantage [46].
Acknowledgments.– The author thanks Frank Ernesto Quintela Rodriguez for pointing out the plasma physics literature and for suggesting the relevance of the present results. Discussions with Vasco Cavina and Marcus Bonança are gratefully acknowledged. The material presented here is based on lectures delivered at the University of Pisa, Italy, in the academic year 2024/25.
I Appendix
I.1 Equivalence of Eq. (6) and (20)
Equation (19) holds regardless of the specific choice of decreasing function $f$. Choosing $f=R$, with $R$ in Eq. (1) we get $P_{1}(\Omega)=R^{-1}(E_{1}(\Omega)$). Furthermore with the specific choice $f=R$, the function $\Omega_{1}$ boils down to the identity. In fact we have
$\Omega_{1}(x)$ $=\int d\mathbf{z}\theta(x-R(\rho_{0}(\mathbf{z})))$
$=\int d\mathbf{z}\theta(\rho_{0}(\mathbf{z})-R^{-1}(x))=R(R^{-1}(x))=x\,.$ (36)
Therefore the inverse of $\Omega_{1}$ is also the identity, i.e., $E_{1}(\Omega)=\Omega$, in this specific case. Summing up
$P_{1}(\Omega)=R^{-1}(\Omega)\,$, (37)
showing that Eq. (20) coincides with Eq. (6).
I.2 Proof of Eqs. (27)
We first note that, for any function $f$, it is
$\int d\mathbf{z}\,\rho(\mathbf{z})f(H_{0}(\mathbf{z}))=\int d\mathbf{z}\mathcal{D}[\rho](\mathbf{z}))f(H_{0}(\mathbf{z}))\,.$ (38)
In fact we have:
$\int d\mathbf{z}\mathcal{D}[\rho](\mathbf{z})f(H_{0}(\mathbf{z}))$
$=\int d\mathbf{z}\int d\mathbf{z}^{\prime}\,\rho(\mathbf{z})\delta[\Omega_{0}(H_{0}(\mathbf{z}^{\prime}))-\Omega_{0}(H_{0}(\mathbf{z}))]f(H_{0}(\mathbf{z}))$
$=\int d\mathbf{z}\int de\,\omega_{0}(e)\rho(\mathbf{z})\delta[\Omega_{0}(e)-\Omega_{0}(H_{0}(\mathbf{z}))]f(H_{0}(\mathbf{z}))$
$=\int d\mathbf{z}\int d\Phi\,\rho(\mathbf{z})\delta[\Phi-\Omega_{0}(H_{0}(\mathbf{z}))]f(H_{0}(\mathbf{z}))$
$=\int d\mathbf{z}\,\rho(\mathbf{z})f(H_{0}(\mathbf{z}))\,$,
where in the second line we have used the definition (26), we then did the usual changes of variables $\mathbf{z}^{\prime}\to e\to\Phi$ (see, e.g., Eq. 6) and finally used $\int d\Phi\delta(\Phi-y)=1$.
Note that a special case of Eq. (38) is
$\int d\mathbf{z}\rho(\mathbf{z})H_{0}(\mathbf{z})=\int d\mathbf{z}\mathcal{D}[\rho(\mathbf{z}))H_{0}(\mathbf{z})\,$, (39)
saying that the dephasing operator $\mathcal{D}$ does not alter the energy expectation.
As a consequence of Eq. (38) we have:
$C[\rho]=\mathcal{H}[\mathcal{D}[\rho]]-\mathcal{H}[\rho]\,$, (40)
where $\mathcal{H}$ denotes the classical information
$\mathcal{H}[\sigma]=-\int d\mathbf{z}\,\sigma(\mathbf{z})\ln\sigma(\mathbf{z})\,.$ (41)
To see that note, from Eq. (26), that $\mathcal{D}[\rho]$ depends on $\mathbf{z}$ through $H_{0}$, Eq. (26), namely it is of the form $f(H_{0})$. Therefore
$C[\rho]$ $=D[\rho||\mathcal{D}[\rho]]$
$=\int d\mathbf{z}\rho\ln\rho-\int d\mathbf{z}\,\rho\ln\mathcal{D}[\rho]$
$=\int d\mathbf{z}\rho\ln\rho-\int d\mathbf{z}\,\mathcal{D}[\rho]\ln\mathcal{D}[\rho]\,.$ (42)
Another useful formula is:
$D[\sigma||\rho_{\beta}]=\beta\int d\mathbf{z}(\sigma(\mathbf{z})-\rho_{\beta}(\mathbf{z}))H_{0}(\mathbf{z})-\mathcal{H}[\rho]+\mathcal{H}[\rho_{\beta}]\,$, (43)
where $\rho_{\beta}$ is a thermal state:
$\rho_{\beta}(\mathbf{z})=\frac{e^{-\beta H_{0}(\mathbf{z})}}{Z_{0}},\quad\beta>0\,.$ (44)
The proof goes as follows
$D[\sigma||\rho_{\beta}]$ $=\int d\mathbf{z}\,\sigma\ln\sigma-\int d\mathbf{z}\,\sigma\ln\rho_{\beta}$
$=-\mathcal{H}[\sigma]-\int d\mathbf{z}\,\sigma\ln(-\beta H_{0}-\ln Z_{0})$
$=-\mathcal{H}[\sigma]+\beta\int d\mathbf{z}\,\sigma H_{0}-\beta\int d\mathbf{z}\,\rho_{\beta}H_{0}+\mathcal{H}[\rho_{\beta}]\,$,
where, in the last line we used the well known formula:
$\mathcal{H}[\rho_{\beta}]=\beta\int d\mathbf{z}\,\rho_{\beta}(\mathbf{z})H_{0}(\mathbf{z})+\ln Z_{0}\,.$ (45)
We are now ready to prove Eq. (27). We have:
$\beta\mathcal{E}_{c}^{c}$ $=\beta(\mathcal{E}_{c}-\mathcal{E}_{c}^{i})=\beta\int d\mathbf{z}H_{0}(\mathcal{P}[\mathcal{D}[\rho]]-\mathcal{P}[\rho])$
$=\beta\int d\mathbf{z}H_{0}(\mathcal{P}[\mathcal{D}[\rho]]-\rho_{\beta})-\beta\int d\mathbf{z}H_{0}(\mathcal{P}[\rho]-\rho_{\beta})$
$=D[\mathcal{P}[\mathcal{D}[\rho]]||\rho_{\beta}]+\mathcal{H}[\mathcal{P}[\mathcal{D}[\rho]]]-\mathcal{H}[\rho_{\beta}]$
$\phantom{xxx}-D[\mathcal{P}[\rho]||\rho_{\beta}]-\mathcal{H}[\mathcal{P}[\rho]]+\mathcal{H}[\rho_{\beta}]\,.$
Equation (27) follows by noticing that $\mathcal{H}[\mathcal{P}[\rho]]=\mathcal{H}[\mathcal{\rho}]$, $\mathcal{H}[\mathcal{P}[\mathcal{D}[\rho]]]=\mathcal{H}[\mathcal{D}[\rho]]$, because, by construction, any state (be it $\rho$ or $\mathcal{D}[\rho]$), is linked to its passive companion ($\mathcal{P}[\rho]$ or $\mathcal{P}[\mathcal{D}[\rho]]$, respectively) by a volume preserving map, and it is a well known fact that the classical information is invariant under such transformations.

### v2

@@
-Extraction of classical ergotropy
+Unified theory of classical and quantum ergotropy
 Abstract
-Finding the time dependent perturbation that extracts the maximal amount of energy (a.k.a.
-ergotropy) from a thermally isolated quantum system is a central, solved, problem in quantum thermodynamics.
-Notably, the same problem has been long studied for classical systems as well, e.g., in the field of plasma physics, but a general solution is still missing there.
-By building on the analogy with the quantum solution, we provide the classical ergotropy extraction driving: it consists of an instantaneous quench followed by an adiabatic return.
-We illustrate how the solution, which is valid under an ergodic assumption, is instrumental to finding the ergotropy extracting driving in more general cases.
-We also show that, just like in the quantum case, the classical ergotropy splits into a coherent and an incoherent part.
-The presented results open new ways for practical energy recovery in the classical regime while suggesting that there is nothing genuinely quantum in the quantum ergotropy problem.
+Quantifying the ergotropy (a.k.a.
+available energy), namely the maximal amount of energy that can be extracted from a thermally isolated system, is a central problem in quantum thermodynamics.
+Notably, the same problem has been long studied for classical systems as well, e.g.
+in plasma physics and astrophysics, where the basic principles for its solution are known for the case of collisionless fluids.
+Here we provide the general analytical expression of ergotropy of classical systems valid regardless of their size and the type of interparticle interactions, and show that it emerges as the classical limit of the quantum expression of ergotropy, for quantum systems that are classically ergodic.
+We thus establish a unified theory of classical and quantum ergotropy, whose applicability ranges from atomic to galactic scale.
+Such unified theory is indispensable for studying the genuine quantum signatures of ergotropy: We show that the celebrated decomposition of quantum ergotropy into coherent ant inchoherent parts survives in the classical regime, indicating that coherences do not necessarily reveal quantumness.
+The unified theory also allows to port tools and methods across the classical-quantum boundary to unlock the solution of standing problems.
+We apply this to swiftly solve the open problem of ergotropy extraction in the classical regime.
 What is the maximal amount of energy that can be extracted from a thermally isolated system by means of a cyclical external driving?
 And what is the driving that achieves such maximal energy extraction?
-These questions are in the limelight of current investigation in the field of quantum thermodynamics [1, 2, 3, 4] where they naturally emerge in the study of quantum batteries [5, 6, 7, 8], quantum heat engines [9, 10, 11, 12, 13] and dynamic cooling [14, 15, 16, 17].
-While addressing these questions for quantum systems has a long history [18] they took momentum only with the work of Allahverdyan et al.
-[19], who coined the expression “ergotropy” for the maximal energy extractable from a quantum thermally isolated system, and provided its analytical expression along with the expression of the according "ergotropy extracting" unitary operator, see Eqs.
-(9,13) below.
-Most notably the ergotropy extraction problem has been long studied for classical systems as well.
-For example, in the field of plasma physics, where, e.g., the recovery of the maximal amount of energy from a plasma of fusion products constitutes a pressing practical challenge [20, 21, 22, 23, 24].
+These questions are in the limelight of current investigation in the field of quantum thermodynamics [1, 2, 3, 4] where they naturally emerge in the study of quantum batteries [5, 6, 7, 8, 9], quantum heat engines [10, 11, 12, 13, 14] and dynamic cooling [15, 16, 17, 18, 19].
+The concept of maximal energy extractable from a thermally isolated system was first introduced, in the context of quantum theory, in 1976 in Ref.
+[20] where the expression “adiabatic availability’ was used to denote it.
+However the topic took momentum only later with the work of Allahverdyan et al.
+[21], who coined the expression “ergotropy” for the same quantity, and provided its analytical expression along with the expression of the according "ergotropy extracting" unitary operator, see Eqs.
+(2,6) below.
+Here we note that the very same quantity has been long studied for classical systems as well, where it is known as “available energy” (often indicated by the symbol Æ).
+The study of available energy in the classical realm can be traced back to Lorenz who introduced it in atmospheric physics in 1955 [22] and has found applications ranging from plasma physics, [23, 24, 25, 26, 27, 28, 29], to astrophysics [30, 31] and climate science [32].
+In this context, available energy is a fundamental tool for the study of instabilities: the higher the available energy the farther the system is from a stable equilibrium.
 In the pioneering and elegantly succinct work of Ref.
-[25] Gardner pinpointed the fundamental principles that guide the calculation of the maximal energy extraction from a classical system by means of phase space volume preserving maps, a.k.a.
-"Gardner free energy" [20, 21].
-As pointed out recently [23] any volume preserving map can be arbitrarily well approximated by a symplectic map, therefore, formally, the Gardner free energy is indeed the maximal energy that can be extracted by a Hamiltonian flow, i.e., the classical ergotropy.
-Therefore, as of today, determining the classical ergotropy is a problem with a known solution.
-However, the pressing practical question of what is the classical ergotropy extraction driving remains an open question.
-Here we first re-derive the classical ergotropy and express it in a new way, Eq.
-(7), and then demonstrate a general driving protocol that achieves it.
-The protocol is intuituve, physically motivated and is inspired by the according solution of the quantum problem whose close analogy with the classical problem is demonstrated.
-The presented protocol works under the provision of the ergodic hypothesis, however we shall discuss how it in fact can be employed as the basis for designing the ergotropy extraction drivings in more general cases as well.
-This opens new ways for the practical implementation of energy extraction protocols in the classical regime.
-The presented theory also sheds further light on the quantum ergotropy, showing that there is very little (if not nothing at all) genuinely quantum in it, besides the discreteness of variables.
-Gardner free energy (a.k.a.
-classical ergotropy).– Given a classical system with unperturbed Hamiltonian $H_{0}(\mathbf{z})$ being initially in a statistical state described by the phase space distribution $\rho_{0}(\mathbf{z})$, Gardner asked what is the distribution $\rho_{1}(\mathbf{z})$ featuring the smallest energy expectation among all those that can be connected to $\rho_{0}$ by a phase-volume preserving map [25].
-Gardner concluded that (i) $\rho_{1}(\mathbf{z})$ must be a decreasing function $g$ of the system unperturbed Hamiltonian $H_{0}(\mathbf{z})$, (ii) for any positive real number $\sigma$, the volume of phase space where $\rho_{1}>\sigma$ must be equal to that where $\rho_{0}>\sigma$  [^1] A simple aftermath of Gardner’s prescription is that one cannot extract energy from states of the form $\rho_{0}=g(H_{0})$ with $g$ a decreasing fuction, a fact that has been formally proved also in Ref.
-[27], see also [28] for a more accessible proof..
-These prescriptions form the basis of the so-called "Gardner restacking algorithm" and allow to calculate the "Gardner free energy" and the function $g$ that determines the “Gardner ground state” $\rho_{1}=g(H_{0})$ [20, 21], a.k.a.
-the “passive state” relative to $H_{0}$ [27, 28].
-Let
-$R(\sigma)=\int d\mathbf{z}\theta[\rho_{0}(\mathbf{z})-\sigma]$ (1)
-denote the measure of phase space space where $\rho_{0}>\sigma$ (the symbol $\theta[\cdot]$ denotes Heaviside step function). $R$ is a decreasing, hence invertible, function. Let
-$\Omega_{0}(E)=\int d\mathbf{z}\theta[E-H_{0}(\mathbf{z})]$ (2)
-[…]
-$R(\sigma)=\int d\mathbf{z}\theta[g(H_{0}(\mathbf{z}))-\sigma]\,$, (3)
-or, equivalently,
-$R(\sigma)$ $=\int d\mathbf{z}\theta[g^{-1}(\sigma)-H_{0}(\mathbf{z})]=\Omega_{0}(g^{-1}(\sigma))\,.$ (4)
-Applying $E_{0}$ to both sides of the above equation, we get
-$E_{0}(R(\sigma))=g^{-1}(\sigma)\,.$ (5)
-The Gardner ground state is therefore $\rho_{1}(\mathbf{z})=g(H_{0}(\mathbf{z}))$ with $g$ obtained by inverting Eq.
-(5)  [^2] Note that since $E_{0}$ is increasing, and $R$ is decreasing, $g^{-1}$ is decreasing, therefore $g$ is decreasing as expected..
-The minimal energy reachable by volume preserving maps reads, accordingly:
-$\breve{E}_{c}$ $=\int d\mathbf{z}H_{0}(\mathbf{z})g(H_{0}(\mathbf{z}))=\int_{0}^{\infty}de\omega_{0}(e)eg(e)$
-$=$ $\int_{0}^{\infty}d\Phi E_{0}(\Phi)g(E_{0}(\Phi))=\int_{0}^{\infty}d\Phi E_{0}(\Phi)R^{-1}(\Phi)\,$, (6)
-where in the first line we changed the integration variable from phase-space $\mathbf{z}$ to energy $e$, using the density of states $\omega_{0}(e)=\Omega_{0}^{\prime}(e)$, in the second line we first made the change of variable $e\rightarrow\Phi=\Omega_{0}(e)$, and then we used Eq.
+[23] Gardner pinpointed the fundamental principles that guide the calculation of the available energy of a collisionless fluid obeying Vlasov dynamics.
+Their application ranges from plasmas [24, 25] to self-gravitating matter [30].
+Here we note that Gardner’s principles can be applied to the calculation of ergotropy of any classical (possibly many-body interacting) Hamiltonian system, whose phase space density evolution is Liouvillean.
+We show how they univocally lead to the general analytical expression of ergotropy of a classical system, see Eqs.
+(15,22).
+Our main result is that, for quantum systems that are classically ergodic, such expression emerges as the semi-classical limit of the celebrated quantum expression of ergotropy [21], Eq.
 (6).
-Remarkably, by inverting the functions $R,\Omega_{0}$ in Eqs.
-(1,2) one swiftly gets $\breve{E}_{c}$, hence the Gardner free energy, that is, the classical ergotropy,
-$\mathcal{E}_{c}=\int d\mathbf{z}H_{0}(\mathbf{z})\rho_{0}(\mathbf{z})-\int_{0}^{\infty}d\Phi E_{0}(\Phi)R^{-1}(\Phi)\,.$ (7)
-Despite the interest of this and similar equivalent expressions [20, 21], they tell nothing about the pressing practical question of how to engineer a time dependent driving $V(\mathbf{z},t)$ such that the system state $\rho_{0}$ will indeed evolve onto the “Gardner ground state” $\rho_{1}$ (a.k.a the passive state of $\rho_{0}$) under the flow generated by the total Hamiltonian $H(\mathbf{z},t)=H_{0}(\mathbf{z})+V(\mathbf{z},t)$ in some time span $0,\tau]$.
-The Quench-Adiabat protocol.– To get a handle on the above issue let’s look at the analogous quantum problem.
+Thus we bridge and unify results obtained in distant subfields of physical investigation, such as plasma- and astro-phyiscs and quantum thermodynamics, while greatly extending their realm of applicability.
+The present unified theory of ergotropy provides the solid theoretical ground for unlocking open questions both in the classical and quantum realms.
+For example a timely topic of quantum thermodynamic investigation concerns the identification of the genuinely quantum features of non-equilibrium energy exchange processes [1, 2, 3, 4].
+Current expectation, based on the recently reported decomposition of quantum ergotropy into coherent and inchoherent parts [33], is that coherences reveal its genuinely quantum features.
+Here we prove that the decomposition survives in the classical regime thus showing that coherences do not necessarily reveal quantumness.
+Clearly, a unified theory also allows to port ideas and methods across the quantum-classical boundary.
+This can be used as a strategy to address open problems on both sides.
+For example we apply it to solving the pressing, yet unsolved, problem of finding a driving protocol that extracts the ergotropy in the classical case.
+Quick review of ergotropy of a quantum system
 Given a quantum system with Hamiltonian $\hat{H}_{0}$ being in a state described by the density operator $\hat{\rho}_{0}$, we ask what is the unitary evolution operator $\hat{U}$ that leads to the state $\hat{\rho}_{1}=\hat{U}\hat{\rho}_{0}\hat{U}^{\dagger}$ of lowest energy expectation
@@
 For Hamiltonians with discrete spectrum acting on a Hilbert space of finite dimension $d$ the above minimum is reached by the unitary operator [21]
-$\hat{U}=\sum_{k=1}^{d}|e_{k}\rangle\langle r_{k}|\,$, (9)
+$\hat{U}=\sum_{k=1}^{d}e^{-i\phi_{k}}|e_{k}\rangle\langle r_{k}|\,$, (2)
 […]
@@
 where $\delta_{jk}$ denotes the Kronecker delta.
-Note that the optimal unitary $\hat{U}$ is exactly realising the quantum analogue of Gardner prescriptions [25]: the “ground state” $\hat{\rho}_{1}$, Eq.
-(12) is a decreasing function of $\hat{H}_{0}$, and the number of eigenstates of initial and final density operators ($\hat{\rho}_{0}$ and $\hat{\rho}_{1}$), whose eigenvalue is less than some value $\sigma$ are the same, for any $\sigma>0$.
-This is so because the density operator eigenvalues are invariant under unitary evolution.
-In classical mechanics, the analogous condition is dictated by the fact that the phase space density is invariant under Liouville evolution.
-What time-dependent perturbation $\hat{V}(t)$ should one enact so that the according Hamiltonian
-$\hat{H}(t)=\hat{H}_{0}+\hat{V}(t)$ (14)
-would induce the dynamical evolution $\hat{U}$?
-Notably the answer is not unique.
-From the mathematical point of view, perhaps the simplest solution is to chose $\hat{H}(t)=\hat{K}\doteq\hbar\frac{i}{\tau}\log U$, for $t\in[0,\tau]$, so that trivially one has $e^{-i\hat{K}\tau/\hbar}=\hat{U}$ in the time span $0,\tau]$.
-From the physical viewpoint, we note that $\hat{U}$ can be implemented by the following quench-adiabat (QA) protocol: (i) instantaneously quench at $t=0$ to
-$\hat{H}_{1}=f(\hat{\rho}_{0})$ (15)
+Ergotropy of a classical system
+Consider a classical system with unperturbed Hamiltonian $H_{0}(\mathbf{z})$ being initially in a statistical state described by the phase space distribution $\rho_{0}(\mathbf{z})$.
+Here $\mathbf{z}=(\mathbf{p},\mathbf{q})$ denotes a point in phase space.
+We ask what is the distribution $\rho_{1}(\mathbf{z})$ featuring the smallest energy expectation among all those that can be connected to $\rho_{0}$ by a Hamiltonian flow.
+Two ingredients are sufficient to answer this question.
+First, as as formally demonstrated in [34], each and all states of the form $\rho_{0}=g(H_{0})$ with $g$ a non-increasing function are “passive” states, meaning are such that no energy can be further extracted therefrom.
+This means that $\rho_{1}$ must be of the form:
+$\rho_{1}(\mathbf{z})=g(H_{0}(\mathbf{z})),\quad g(x)\leq g(y),\text{for }x>y$ (7)
+Second, not only any Hamiltonian flow preserves the volume in phase space (Liouville theorem) but also, as noted in [27], any volume preserving map can be arbitrarily well approximated by some Hamiltonian flow.
+This extends the optimisation space from the set of Hamiltonian maps, to the more easily treatable set of “volume preserving" maps.
+In particular this would imply that for any $r>0$, the volume of phase space where $\rho_{1}>r$ must be equal to that where $\rho_{0}>r$, that is:
+$\int d\mathbf{z}\theta(\rho_{0}(\mathbf{z})-r)=\int d\mathbf{z}\theta(\rho_{1}(\mathbf{z})-r)$ (8)
+where $\theta$ denotes Heaviside step function.
+Eqs.
+(7,8) where first noted by Gardner [23] for the case of a collisionless plasma, for which $\rho_{0}$ is a single particle distribution obeying Vlasov equation.
+Their validity is however general: they hold for any (possibly many-body interacting) Hamiltonian system.
+Most importantly they univocally single-out the function $g$ that determines the passive companion, $\rho_{1}$, of $\rho_{0}$.
+To see that, let
+$\Sigma(r)=\int d\mathbf{z}\theta[\rho_{0}(\mathbf{z})-r]$ (9)
+denote the measure of phase space where $\rho_{0}>r$ (the symbol $\theta[\cdot]$ denotes Heaviside step function).
+$\Sigma$ is a decreasing, hence invertible, function.
+We shall soon comment on the physical meaning of the function $\Sigma^{-1}$.
+Let
+$\Omega_{0}(E)=\int d\mathbf{z}\theta[E-H_{0}(\mathbf{z})]$ (10)
+[…]
+$\Sigma(r)$ $=\int d\mathbf{z}\theta[g(H_{0}(\mathbf{z}))-r]$ (11)
+$=\int d\mathbf{z}\theta[g^{-1}(r)-H_{0}(\mathbf{z})]=\Omega_{0}(g^{-1}(r))\,$, (12)
+that is $\Sigma=\Omega_{0}\circ g^{-1}$.
+It follows that $g=\Sigma^{-1}\circ\Omega_{0}$  [^1] Note that since $\Sigma^{-1}$ is decreasing, and $\Omega_{0}$ is increasing, $g$ is decreasing as expected..
+Combining with Eq.
+(7) we get:
+$\rho_{1}(\mathbf{z})=\Sigma^{-1}(\Omega_{0}(H_{0}(\mathbf{z})))\,.$ (13)
+This explicit expression of the passive companion, $\rho_{1}$, of $\rho_{0}$, in Eq.
+(13) is crucial for what follows.
+It generalises similar ones obtained in plasma/astro-physics for collisionless systems [30, 25].
+Among plasma physicists $\rho_{1}$ is known as the “ground state associated to $\rho_{0}$", while astrophysicists call it the “equilibrium associated to $\rho_{0}$" [30].
+Given $\rho_{1}$, Eq. (13), the minimal energy reachable by volume preserving maps reads:
+$\breve{E}_{c}$ $=\int d\mathbf{z}H_{0}(\mathbf{z})\rho_{1}(\mathbf{z})=\int d\mathbf{z}H_{0}(\mathbf{z})\Sigma^{-1}(\Omega_{0}(H_{0}(\mathbf{z})))$
+$=\int_{0}^{\infty}de\,\omega_{0}(e)e\Sigma^{-1}(\Omega_{0}(e))=\int_{0}^{\infty}d\Phi E_{0}(\Phi)\Sigma^{-1}(\Phi)$ (14)
+where we first changed the integration variable from phase-space $\mathbf{z}$ to energy $e=H_{0}(\mathbf{z})$, using the density of states $\omega_{0}(e)=\Omega_{0}^{\prime}(e)$, and then made the change of variable $e\rightarrow\Phi=\Omega_{0}(e)$.
+The ergotropy $\mathcal{E}_{c}$ (aka available energy) of a classical system is then:
+$\mathcal{E}_{c}=\int d\mathbf{z}H_{0}(\mathbf{z})\rho_{0}(\mathbf{z})-\int_{0}^{\infty}d\Phi E_{0}(\Phi)\Sigma^{-1}(\Phi)\,.$ (15)
+Note that since $E_{0}$ has the units of energy, $\Sigma^{-1}$ has units of a probability density in “phase volume”.
+In fact it can be interpreted as the probability density, call it $P_{1}(\Phi)$, that the system is found on the eigensurface of $H_{0}$ that encloses the volume $\Phi$ provided it is distributed according to $\rho_{1}$.
+To see that we express $P_{1}$ as the overlap of $\rho_{1}$ and the microcanonical state $\rho_{\mu}$ that “lives” on that eingensurface
+$\rho_{\mu}(\mathbf{z};\Phi)=\delta[\Phi-\Omega_{0}(H_{0}(\mathbf{z}))]$ (16)
+Note that here the microcanonical distribution is parametrised via the enclosed-volume $\Phi$ rather than the usual energy $E$, which makes it naturally normalised.
+Such unusual, yet convenient, parametrisation, which is crucial in this work, was first put forward in Ref.
+[43].
+As anticipated above, we have:
+$P_{1}(\Phi)$ $=\int d\mathbf{z}\rho_{1}(\mathbf{z})\delta[\Phi-\Omega_{0}(H_{0}(\mathbf{z}))]$ (17)
+$=\int d\mathbf{z}$ $\Sigma^{-1}(\Omega_{0}(H_{0}(\mathbf{z})))\delta[\Phi-\Omega_{0}(H_{0}(\mathbf{z}))]=\Sigma^{-1}(\Phi)\,.$
+Introducing the probability density $P_{0}(\Phi)$ of finding the system on the eigensurface of $H_{0}$ that encloses the volume $\Phi$, provided it is distributed according to $\rho_{0}$
+$P_{0}(\Phi)=\int d\mathbf{z}\rho_{0}(\mathbf{z})\delta[\Phi-\Omega_{0}(H_{0}(\mathbf{z}))]\,.$ (18)
+the ergotropy takes the symmetric and meaningful form:
+$\mathcal{E}_{c}=\int d\Phi[P_{0}(\Phi)-P_{1}(\Phi)]E_{0}(\Phi)\,.$ (19)
+Most remarkably $P_{0}(\Phi)$ is linked to the probability density $P_{1}$ as follows:
+$P_{0}(\Phi)$
+$=\int d\Theta\int d\mathbf{z}\delta[\Phi-\Omega_{0}(H_{0}(\mathbf{z}))]\delta[\Theta-\Sigma(\rho_{0}(\mathbf{z}))]\rho_{0}(\mathbf{z})$
+$=\int d\Theta\int d\mathbf{z}\delta[\Phi-\Omega_{0}(H_{0}(\mathbf{z}))]\delta[\Theta-\Sigma(\rho_{0}(\mathbf{z}))]\Sigma^{-1}(\Theta)$
+$=\int d\Theta G[\Theta|\Phi]P_{1}(\Theta)\,$, (20)
+where we first inserted the unit resolution, $1=\int d\Theta\delta(\Theta-y)$, and then introduced the overlap
+$\hskip-1.42262ptG[\Theta|\Phi]=\int d\mathbf{z}\delta[\Phi-\Omega_{0}(H_{0}(\mathbf{z}))]\delta[\Theta-\Sigma(\rho_{0}(\mathbf{z}))]$ (21)
+between the microcanonical states having as support the eigensurfaces of $\rho_{0}$ and $H_{0}$ enclosing the volumes $\Theta$ and $\Phi$, respectively [43]. Combining Eqs. (19, 20) we finally get
+$\mathcal{E}_{c}=\int d\Phi d\Theta P_{1}(\Theta)(G[\Theta|\Phi]-\delta[\Phi-\Theta])E_{0}(\Phi)\,.$ (22)
+Note the formal analogy with the quantum formula (6) whereby the discrete principal quantum numbers $k,j$ are replaced by the continuous "enclosed volumes" variables $\Theta,\Phi$, and the quantum overlap of energy eigenstates is replaced by the classical overlaps of microcanonical states.
+The overlap $G$ can be understood as a special case of a classical propagator of a phase-volume density, which is the analogous of a quantum transition matrix.
+Such propagators were first discussed in Refs.
+[45, 46]
+Semiclassical limit
+The analogy between Eqs.
+(6, 22) is not just aesthetic, but it is substantial.
+In order to see that consider the phase space formulation of quantum mechanics, based on the Wigner-Weyl transform [47, 48].
+Note first of all that the Weyl representation $H^{W}_{0}(\mathbf{z})$ of the operator $\hat{H}_{0}$ tends in the semi-classical limit to its classical expressions $H_{0}(\mathbf{z})$, where the position and momentum operators are replaced by real variables.
+Next note that the Wigner function $W_{\hat{\rho}_{0}}(\mathbf{z})$ of the initial preparation tends to a proper Liouville density $\rho_{0}(\mathbf{z})$; similarly, the final state $W_{\hat{\rho}_{1}}(\mathbf{z})$ tends to the Liouville density $\rho_{1}(\mathbf{z})$.
+Also the latter is linked to the former by the Liouville dynamics because the Moyal dynamics dictating the evolution of the Wigner functions becomes Liouvillean in the classical limit [47].
+Let’s then consider the probability $r_{k}$. According to Eq. (5) it is
+[…]
+[…]
+[…]
+go over the continuous function
+$E_{0}(\Phi)=\int d\mathbf{z}H_{0}(\mathbf{z})\delta(\Phi-\Omega_{0}(H_{0}(\mathbf{z})))$ (25)
+specifying the energy of each microcanonical eigenstate, labelled by the phase volume enclosed by their support.
+Finally, the quantum transition probabilities can be expressed as:
+[…]
+Assuming that the Hamiltonian $\hat{H}_{1}=f(\hat{\rho}_{0})$ (with $f$ a generic decreasing function), is also classically ergodic, they go in the semi-classical limit, over the classical overlaps $G[\Theta|\Omega]$, Eq.
+(21).
+See also Ref.
+[51] for explicit examples.
+All these observations demonstrate that, under the provision of the ergodic hypothesis, the classical ergotropy, Eq.
+(15) emerges as the classical limit of the quantum ergotropy, Eq.
+(6).
+Extraction of ergotropy
+The above theory allows to apply methods and tools developed in the quantum case to the classical case and vice-versa.
+As an example we consider the problem of ergotropy extraction of a classical system.
+Notably, despite its practical interest, this problem has not been addressed so far in the literature.
 […]
@@
 $H_{1}(\mathbf{z})=f(\rho_{0}(\mathbf{z}))$ (27)
-[…]
-To prove that the QA protocol extracts the calssical ergotropy let
-$P_{1}(\Omega)$ $=\int d\mathbf{z}\rho_{0}(\mathbf{z})\delta[\Omega-\Omega_{1}(H_{1}(\mathbf{z}))]$ (17)
-denote the probability density of finding, at $t=0$, the system on the iso-$H_{1}$ hyper-surface that encloses the volume of phase space $\Omega$. The symbol $\delta$ denotes Dirac’s delta and
-$\Omega_{1}(E)=\int d\mathbf{z}\theta(E-H_{1}(\mathbf{z}))$ (18)
-denotes the volume of phase space enclosed by the hypersurface $H_{1}=E$. We have
-$P_{1}(\Omega)$ $=\int d\mathbf{z}f^{-1}(H_{1}(\mathbf{z}))\delta[\Omega-\Omega_{1}(H_{1}(\mathbf{z}))]$
-$=\int de\omega_{1}(e)f^{-1}(e)\delta[\Omega-\Omega_{1}(e)]$
-$=\int d\Phi f^{-1}(E_{1}(\Phi))\delta[\Omega-\Phi]=f^{-1}(E_{1}(\Omega)),$ (19)
-where in the first line we used $\rho_{0}=f^{-1}(H_{1})$, in the second line we changed the integration variable from phase-space $\mathbf{z}$ to energy $e$, using the density of states $\omega_{1}(e)=\Omega_{1}^{\prime}(e)$, and in the third line we made the change of variable $e\rightarrow\Phi=\Omega_{1}(e)$, with $E_{1}=\Omega_{1}^{-1}$ denoting its inverse.
-Under the ergodic hypothesis, because of adiabatic invariance of the phase volume, the quantity $P_{1}(\Omega)$ is in fact equal to the probability density of finding the system on a hypersurface of constant $H_{0}$, enclosing a volume $\Omega$, at the end of the adiabatic driving [42, 28].
-Therefore the final mean energy reads:
-$\breve{E}_{c}=\int d\Omega P_{1}(\Omega)E_{0}(\Omega)\,.$ (20)
-As shown in the appendix, the above expression is in fact identical to Gardner ground state energy Eq. (7), thus concluding the proof.
-Quantum-classical analogy.– The classical ergotropy amounts to $\mathcal{E}_{c}=\int d\mathbf{z}\rho(\mathbf{z})H_{0}(\mathbf{z})-\int d\Omega P_{1}(\Omega)E_{0}(\Omega)$, or, in more symmetric form:
-$\mathcal{E}_{c}=\int d\Omega[P_{0}(\Omega)-P_{1}(\Omega)]E_{0}(\Omega)\,$, (21)
-where
-$P_{0}(\Omega)=\int d\mathbf{z}\delta[\Omega-\Omega_{0}(H_{0}(\mathbf{z}))]\rho_{0}(\mathbf{z})$ (22)
-[…]
-$P_{0}(\Omega)$
-$=\int d\Theta\int d\mathbf{z}\delta[\Omega-\Omega_{0}(H_{0}(\mathbf{z}))]\delta[\Theta-\Omega_{1}(H_{1}(\mathbf{z}))]f^{-1}(H_{1}(\mathbf{z}))$
-$=\int d\Theta\int d\mathbf{z}\delta[\Omega-\Omega_{0}(H_{0}(\mathbf{z}))]\delta[\Theta-\Omega_{1}(H_{1}(\mathbf{z}))]f^{-1}(E_{1}(\Theta))$
-$=\int d\Theta G[\Theta|\Omega]P_{1}(\Theta)\,$, (23)
-where we first inserted the unit resolution, $1=\int d\Theta\delta(\Theta-y)$, and then we introduced the “overlap” between the microcanonical states of fixed $H_{0}(\mathbf{z})$ and $H_{1}(\mathbf{z})$, whose support respectively enclose the volumes $\Theta$ and $\Omega$  [^3] When parametrised via the enclosed volume variable $\Phi$, instead of the usual energy $E$, the microcanonical ensemble is naturally normalised and simply reads: $\rho(\mathbf{z};\Phi)=\delta(\Phi-\Omega(H(\mathbf{z})))$, where $\Omega(x)$ is the phase space volume of the region $H\leq x$ [28].
-$G[\Theta|\Phi]$ can be interpreted as the conditional probability that the system is found on the hypersurface of constant $H_{0}$ that encloses the volume $\Omega$ given that it is homogeneously (i.e., microcanonically) distributed on the hypersurface of constant $H_{1}$ that encloses the volume $\Theta$, or viceversa.
-It is a special instance of the propagator $G_{0,\tau}[\Theta|\Omega]=\int d\mathbf{z}\delta[\Omega-\Omega_{0}(H_{0}(\mathbf{z}))]\delta[\Theta-\Omega_{1}(H_{1}(\mathbf{z}_{\tau}))]$ valid for a generic time dependent protocol taking $H_{0}$ to $H_{1}$ in a finite time $\tau$, where $\mathbf{z}_{\tau}$ denotes the evolved of initial condition $\mathbf{z}$ according to the protocol [41, 42, 28].
-For a sudden quench it is $\tau=0$, so that $\mathbf{z}_{\tau}=\mathbf{z}$ and the expression boils down to Eq.
-(24).
-Furthermore just like the quantum transition probabilities $p_{ij}$ form a doubly stochastic matrix, similarly the propagator $G$ is doubly stochastic in the sense that $\int d\Theta G[\Theta|\Omega]=\int d\Omega G[\Theta|\Omega]=1$ [42, 28]
-$G[\Theta|\Omega]=\int d\mathbf{z}\delta[\Omega-\Omega_{0}(H_{0}(\mathbf{z}))]\delta[\Theta-\Omega_{1}(H_{1}(\mathbf{z}))]$ (24)
-Summing up, the classical ergotropy reads:
-$\mathcal{E}=\int d\Omega d\Phi P_{1}(\Theta)(G[\Theta|\Omega]-\delta[\Omega-\Theta])E_{0}(\Omega)\,.$ (25)
-Note the formal analogy with the quantum formula (13) whereby the discrete (adiabatically invariant) principal quantum numbers $k,j$ are replaced by the continuous (adiabatically invariant) "enclosed volumes" variables $\Theta,\Omega$, and the quantum overlap $p_{kj}=|\langle r_{j}|e_{k}\rangle|^{2}$ is replaced by the classical overlap $G[\Theta|\Omega]$  [^4] The analogy becomes even more evident if one notes that, for systems admitting a phase space description, the quantum transitions $p_{kj}$ can be written as the phase space overlaps between the Wigner functions of the eigenstates of $\hat{H}_{0}$ and $\hat{H}_{1}$ [49] and that (for ergodic systems), such Wigner functions approach the classical microcanonical distributions in the classical limit..
-Quite remarkably, just like the quantum ergotropy, the classical ergotropy splits into a coherent and an incoherent part [45]. Let
-$\mathcal{D}[\rho](\mathbf{z})=\int d\mathbf{z}^{\prime}\,\rho(\mathbf{z}^{\prime})\delta[\Omega_{0}(H_{0}(\mathbf{z}^{\prime}))-\Omega_{0}(H_{0}(\mathbf{z}))]$ (26)
-[…]
-$\beta\mathcal{E}_{c}^{c}[\rho]$ $=C[\rho]+D[\mathcal{P}[\mathcal{D}[\rho]||\rho_{\beta}]-D[\mathcal{P}[\rho]||\rho_{\beta}]\,$, (27)
-where
-$\mathcal{E}_{c}^{c}[\rho]\doteq\mathcal{E}_{c}[\rho]-\mathcal{E}_{c}^{i}[\rho]\quad\mathcal{E}_{c}^{i}[\rho]\doteq\mathcal{E}_{c}[\mathcal{D}[\rho]]$ (28)
-define the coherent and incoherent parts of $\mathcal{E}_{c}$, $\mathcal{P}[\rho]$ denotes the passive state (i.e., the Gardner ground state) associated to $\rho$, $\rho_{\beta}=e^{-\beta H_{0}}/Z$ is a thermal state, $D[\cdot||\cdot]$ denotes the Kullback-Leibler divergence, and
-$C[\rho]=D[\rho||\mathcal{D}[\rho]]$ (29)
-quantifies the “coherence” of $\rho$, that is how much it is distinguisgable from its dephased companion $\mathcal{D}[\rho]$.
-Equations (27-29) read exactly like their quantum counterparts where $D$ denotes the quantum Kullback-Leibler divergence, and $\mathcal{D}$ is the quantum dephasing operator relative to $\hat{H}_{0}$ [45].
-Example.–
+for some monotonously decreasing function $f$, (ii) adiabatically return to $H_{0}(\mathbf{z})$.
+Assuming that the motion induced by the “frozen” Hamiltonians $H(\mathbf{z},t)$ is ergodic on their eigensurfaces [39, 43] for all $t$’s in the QA protocol time span, the phase volume $\Omega_{0}$ is an adiabatic invariant [37, 53, 54, 55].
+Consequently the QA protocol would dynamically map the eigensurfaces of $H_{1}(\mathbf{z})$ (that is the eigensurfaces of $\rho_{0}(\mathbf{z})$) onto the eigensurfaces of $H_{0}(\mathbf{z})$ that enclose same phase volumes thus realising the Gardner prescriptions.
+This can be formally proved by noticing that, under the provision of the ergodic adiabatic theorem [37, 53, 54, 55], the propagator of densities in enclosed-volume space is a Dirac delta [45] 222this is analogous to the quantum case where the transition matrix becomes the identity in the adiabatic limit [52].
+Accordingly, the probability density of finding the system on the hypersurface of $H_{0}$ enclosing the volume $\Phi$, at the end of the adiabatic driving would be identical to the probability density $P_{1}(\Phi)$ of finding it initially on the hypersurface of $H_{1}$ (i.e., of $\rho_{0}$) enclosing the same volume.
+The energy change across the QA protocol would then be given by Eq.
+(19), proving that it extracts the ergotropy.
+On a more intuitive level note that, just like in the quantum scenario, with the first step you lock the system in a passive state relative to $H_{1}$, and with the second step you adiabatically map it onto the passive state relative to $H_{0}$.
+This is well illustrated by the example below.
 Figure 1: Panels a-d): Phase space sketch of extraction of ergotropy from a non-stationary Gaussian state, Eq.
-(31) of a harmonic oscillator, Eq.
-(30).
+(28) of a harmonic oscillator.
 a) The phase density rotates under the unperturbed dynamics.
@@
 c,d) In the time span $0,\tau\gg 1]$ the potential is slowly returned to its initial position.
-Panels e-f): Sketches of the states $\rho_{0}$, $\rho^{\prime}$, Eqs.
-(34,35).
-In all panels darker blue denotes higher density.
-The solid lines denote iso-$H_{0}$ curves in phase space
-Consider a 1D Harmonic oscillator
-$H_{0}(q,p)=\frac{p^{2}}{2}+\frac{q^{2}}{2}$ (30)
-being at $t=0$ in the Gaussian state
+Consider a 1D Harmonic oscillator, $H_{0}(q,p)={p^{2}}/{2}+{q^{2}}/{2}$, being at $t=0$ in the Gaussian state
 $\rho_{0}(q,p)=\frac{1}{2\pi\sigma}\exp\left[-\frac{(q-q_{0})^{2}+p^{2}}{2\sigma}\right]\,.$ (28)
@@
 By suddenly displacing the harmonic potential minimum at $q=q_{0}$ blocks the state at the bottom of the displaced harmonic potential: the state is now stationary (in fact passive) relative to the new Hamiltonian
-$H_{1}(q,p)=\frac{p^{2}}{2}+\frac{(q-q_{0})^{2}}{2}=f(\rho_{0}(\mathbf{z}))\,$, (32)
+$H_{1}(q,p)=\frac{p^{2}}{2}+\frac{(q-q_{0})^{2}}{2}=f(\rho_{0}(q,p))\,$, (29)
 where $f(x)=-\sigma\ln(2\pi\sigma x)$ is a decreasing function, see Fig. 1b). By adiabatically moving the harmonic potential minimum back to the origin of the $q$-axis, the state evolves onto
@@
 because each instantaneous Hamiltonian visited by the protocol is ergodic, see Fig. 1c,d). Since the state is passive relative to $H_{0}$ no more energy can be extracted therefrom.
-Discussion.– There are a number of issues that limit the applicability of the QA protocol for ergotropy extraction.
-Let’s consider, for example, an initial state of the harmonic oscillator featuring a central valley, e.g.:
-$\rho_{0}(q,p)$ $=\exp{\left[-\frac{(p^{2}+(q-q_{0})^{2})^{2}}{4}+\frac{p^{2}+(q-q_{0})^{2}}{2}\right]}/N\,$, (34)
-where $N$ is the normalisation factor, see the sketch in Fig.
-1e).
-Any choice of decrasing function $f$ would result in a Mexican hat shaped Hamiltonian $H_{1}$ which could be very hard to be implemented.
-Furthermore since any such $H_{1}$ would feature a central peak, not all regions $H_{1}<E$ would be simply connected, unlike those for which $H_{0}<E$, therefore they cannot be smoothly and adiabatically linked to each other.
-How to get around such problems?
-In this case one could still apply the QA protocol, but with the physical $H_{1}$ in Eq.
-(32), which has the same level hypersurfaces as $\rho_{0}$.
-Having $H_{1}$ the same topology of phase volumes as $H_{0}$, by adiabatically return to $H_{0}$, the state would be mapped onto
-$\rho^{\prime}(q,p)$ $=\exp{\left[-\frac{(p^{2}+q^{2})^{2}}{4}+\frac{p^{2}+q^{2}}{2}\right]}/N=\frac{e^{-H^{2}_{0}+H_{0}}}{N}\,$, (35)
-which is less energetic than $\rho_{0}$, but not yet passive although it is stationary (because it is a function of $H_{0}$, but not a decreasing one).
-In fact $\rho^{\prime}$ has a central valley, see the sketch in Fig.
-1f).
-Using the quantum jargoon, one would say that the state $\rho^{\prime}$ features population inversions.
-In order to decrease the energy further one should permute rings of higher populations with rings of same volume with lower population and higher (average) energy.
-The smaller the rings thickness, the closer one can get to the ergotropy extraction, by increasing the number of such permutations.
-Interestingly, physical drivings that indeed achieve such permutations have been described in Refs.
+Decomopsition of ergotropy in coherent and incoherent parts
+The above unified theory of ergotropy is an indispensable tool for assessing the genuine quantum signatures of ergotropy.
+Ref.
+[33] has highlighted that in the quantum case, the ergotropy splits into a coherent and an incoherent part.
+The standard expectation is that the coherent part reveals the quantum fingerprints of energy extraction in the quantum regime.
+Here we show that, contrary to that expectation, coherent contributions survive the quantum to classical transition, as the decomposition into a coherent and an incoherent part remains unaltered in the classical case.
+The main tool that is needed for this analysis is the classical “dephasing” operator $\mathcal{D}$ relative to the Hamiltonian $H_{0}$
+$\mathcal{D}[\rho](\mathbf{z})=\int d\mathbf{z}^{\prime}\,\rho(\mathbf{z}^{\prime})\delta[\Omega_{0}(H_{0}(\mathbf{z}^{\prime}))-\Omega_{0}(H_{0}(\mathbf{z}))]\,.$ (31)
+$\mathcal{D}$ homogenises a classical distribution $\rho$ over the eigensurfaces of $H_{0}$, thus rendering it “diagonal” with respect to $H_{0}$ (i.e., $\{\mathcal{D}[\rho],H_{0}\}=0$, where $\{\cdot,\cdot\}$ denotes Poisson brakets) while preserving its energy  [^3] The classical dephasing operator was first introduced in Ref.
+[58] where it was expressed in the energy parametrisation.
+Here we express it in the equivalent enclosed-volume representation, which we find more convenient.
+.
+Exactly like in the quantum case [33], we have
+$\mathcal{E}_{c}[\rho]=\mathcal{E}_{c}^{c}[\rho]+\mathcal{E}_{c}^{i}[\rho]$ (32)
+where
+$\mathcal{E}_{c}^{c}[\rho]$ $=\beta^{-1}(C[\rho]+D[\mathcal{P}[\mathcal{D}[\rho]]||\rho_{\beta}]-D[\mathcal{P}[\rho]||\rho_{\beta}])$ (33)
+$\mathcal{E}_{c}^{i}[\rho]$ $=\mathcal{E}_{c}[\mathcal{D}[\rho]]$ (34)
+denote the coherent and incoherent parts, respectively.
+Here, in full analogy with the quantum case $\mathcal{P}[\rho]$ denotes the passive companion of $\rho$, $\rho_{\beta}=e^{-\beta H_{0}}/Z$ is a thermal state, $D[\cdot||\cdot]$ denotes the Kullback-Leibler divergence, and
+$C[\rho]=D[\rho||\mathcal{D}[\rho]]$ (35)
+quantifies the “coherence” of $\rho$, that is how much it is distinguisgable from its dephased companion $\mathcal{D}[\rho]$.
+Equations (32-35) read exactly like their quantum counterparts where $D$ denotes the quantum Kullback-Leibler divergence, and $\mathcal{D}$ is the quantum dephasing operator relative to $\hat{H}_{0}$ [33].
+The proof proceeds in a way that is fully analogous to the quantum case, see appendix for details.
+Note that both in quantum and classical cases lack of commutation between the state $\rho_{0}$ and the Hamiltonian $H_{0}$, either in the sense of quantum commutators or of Poisson brackets, signals that the state $\rho_{0}$ is not a function of $H_{0}$ hence it is not passive and its ergotropy is not null.
+Such lack of commutation, in both cases is associated to coherences, which in the classical case can be visualised as inhomogeneities of the phase distributions over the energy eigensurfaces [58, 59].
+Discussion
+We have developed a unified theory of classical and quantum ergotropy, a.k.a available energy.
+The theory bridges results obtained in distant parts of the literature, i.e., plasma and astro-physics and quantum thermodynamics, by establishing a common dictionary and unifying them, while greatly extending their realm of applicability.
+The theory applies to any system obeying Liouville dynamics, regardless of the dimension $d$ of its phase space.
+This includes collisionless plasmas ($d=6$), single particle systems in D dimensions (e.g., the harmonic oscillator of the above exameple $d=2$), but also many-body interacting systems ($d=6N$), like interacting gases and self-gravitating matter.
+Due to the semi-classical limit, we can in fact safely state that the present theory applies at scales ranging from atomic to galactic.
+We remark however that our derivation of the classical limit relies on the ergodic hypothesis.
+Further investigations are needed to study the non-ergodic case.
+Thanks to the unified picture we were able to address open questions existing both in the quantum and the classical realms.
+On the classical side, an open pressing question is how to practically realise the time dependent perturbation that does indeed extract the ergotropy.
+By porting the simple quantum solution across the quantum-classical boundary, we swiftly formulate the formal solution of this open problem.
+The solution, the QA protocol, is extremely simple and intuitive in hind-sight, but yet never reported so far.
+We remark that its practical implementation may be greatly hindered, depending on the shape of $\rho_{0}$; e.g., when the eigensurfaces of $H_{0}$ and $\rho_{0}$ are not topologically equivalent, it is impossible to link them smoothly and adiabatically.
+We leave a discussion of these challenging problems to future studies while anticipating that they could be solved by combining the present solution with known phase-space methods such as those that dynamically swap regions of phase space nested one into the other [60].
+Despite these difficulties, our results form the basis for the systematic study of the problem of energy extraction in the classical regime.
+On the quantum side an issue that is in the limelight of timely investigation concerns the quantification of genuinely quantum signatures associated to non-equilibrium energy extraction processes.
+The present unified theory is indispensable for a systematic study of this problem.
+Here, for example, we have highlighted that coherences (which in classical system are revealed as inhomogeneities of the phase distribution [58]) are a source of ergotropy both in the classical and quantum cases.
+Furthermore we showed that the recently reported decomposition of ergotropy into coherent and incoherent components, which was suggested to be revealing of the quantum fingerprints of ergotropy [33], survives in the classical regime.
+This indicates that, contrary to a widespread expectation, energy coherences do not necessarily constitute a genuine quantum thermodynamic advantage, which corroborates the view expressed in Ref.
 [58].
-Notably, such permutation protocols defy the conditions of validity of the adiabatic theorem (because they aim at adiabatically connecting topologically inequivalent regions), but, as numerically demonstrated in Ref.
-[48], they can nonetheless approach the wanted permutation with very high accuracy.
-Thus the breakdown of the ergodic hypothesis does not seem to be such a dramatic problem for ergotropy extraction (as much as it is not dramatic for the mechanical foundations of thermodynamics [35, 28]).
-In general it is expected that the QA protocol works well whenever $H_{0}$ and $\rho_{0}$ have topologically equivalent phase space structure.
-In particular no problems occur when the regions $H_{0}\leq E$ and $\rho_{0}>\sigma$ are all simply connected as is the case in the first example above.
-Conclusions.– By bridging between quantum thermodynamics and classical plasma physics, the present work offers new viewpoints to both fields.
-On one hand it provides a guiding principle for energy extraction from classical systems based on the QA protocol (or variants thereof, e.g., as discussed above); on the other it suggests that the quantum theory of ergotropy has no genuinely quantum feature.
-The two pictures, classical and quantum, are fully one the close analogue of the other.
-Specifically a decomposition into coherent and incoherent part exists just like in the quantum case, corroborating the idea that energy coherences do not constitute a genuine quantum thermodynamic advantage [46].
-Acknowledgments.– The author thanks Frank Ernesto Quintela Rodriguez for pointing out the plasma physics literature and for suggesting the relevance of the present results.
+Most likely, quantum fingerprints should be looked among the genuinely quantum aspects of coherence, e.g., quantum entanglement, as suggested in Ref.
+[61].
+The issue remains open but now we have powerful new tools for further systematic investigations.
+Acknowledgments
+The author thanks Frank Ernesto Quintela Rodriguez for pointing out the plasma physics literature and Vittorio Giovannetti for pointing out Ref.
+[34].
 Discussions with Vasco Cavina and Marcus Bonança are gratefully acknowledged.
@@
 I Appendix
-I.1 Equivalence of Eq. (6) and (20)
-Equation (19) holds regardless of the specific choice of decreasing function $f$.
-Choosing $f=R$, with $R$ in Eq.
-(1) we get $P_{1}(\Omega)=R^{-1}(E_{1}(\Omega)$).
-Furthermore with the specific choice $f=R$, the function $\Omega_{1}$ boils down to the identity.
-In fact we have
-$\Omega_{1}(x)$ $=\int d\mathbf{z}\theta(x-R(\rho_{0}(\mathbf{z})))$
-$=\int d\mathbf{z}\theta(\rho_{0}(\mathbf{z})-R^{-1}(x))=R(R^{-1}(x))=x\,.$ (36)
-Therefore the inverse of $\Omega_{1}$ is also the identity, i.e., $E_{1}(\Omega)=\Omega$, in this specific case. Summing up
-$P_{1}(\Omega)=R^{-1}(\Omega)\,$, (37)
-showing that Eq. (20) coincides with Eq. (6).
-I.2 Proof of Eqs. (27)
-We first note that, for any function $f$, it is
-$\int d\mathbf{z}\,\rho(\mathbf{z})f(H_{0}(\mathbf{z}))=\int d\mathbf{z}\mathcal{D}[\rho](\mathbf{z}))f(H_{0}(\mathbf{z}))\,.$ (38)
+To prove Eq. (32) we first note that, for any function $f$, it is
+$\int d\mathbf{z}\,\rho(\mathbf{z})f(H_{0}(\mathbf{z}))=\int d\mathbf{z}\mathcal{D}[\rho](\mathbf{z})f(H_{0}(\mathbf{z}))\,.$ (36)
 In fact we have:
@@
 $=\int d\mathbf{z}\,\rho(\mathbf{z})f(H_{0}(\mathbf{z}))\,$,
-where in the second line we have used the definition (26), we then did the usual changes of variables $\mathbf{z}^{\prime}\to e\to\Phi$ (see, e.g., Eq.
-6) and finally used $\int d\Phi\delta(\Phi-y)=1$.
+where in the second line we have used the definition (31), we then did the changes of variables $\mathbf{z}^{\prime}\to e\to\Phi$, as, in Eq.
+Ergotropy of a classical system, and finally used $\int d\Phi\delta(\Phi-y)=1$.
 Note that a special case of Eq. (36) is